40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 2024     Entry time: Tue Sep 29 23:43:46 2009
Author: rob 
Type: Update 
Category: SUS 
Subject: ITMY UL OSEM 

We had a redo of elog entry 975 tonight.  The noisy OSEM was fixed by jiggling the rack end of the long cable.  Don't know exactly where--I also poked around the OSEM PD interface board.

In the attached PDF the reference trace is the noisy one.

Attachment 1: ITMYosemBAD.pdf  45 kB  | Hide | Hide all
ELOG V3.1.3-